OpenStudy (anonymous):

a box of matches measures 2.25inx.875inx.375in. a lawn is 165ftx310ft. how many match boxes would fit in the yard at 8 feet deep?

OpenStudy (anonymous):

i tried taking to total area of the field and dividing it by the total area of the boxes but it game me a number that was not a choice

OpenStudy (anonymous):

area of matchbox = 2.25 * 0.875 = 1.96875 sq ins when you say 8 ft deep what exactly do you mean?

OpenStudy (anonymous):

the match box is 2.25 inches long, .875 inches wide, and .375 inches high. the field is 165 feet long and 310 feet wide. how many boxes would it take to fill the field up to 8 feet deep with match boxes?

OpenStudy (anonymous):

ok thanx then you need volume of match box in cu ins, the volume of pile of matches in cu ins and divide first into second - that'll be a lot of boxes!!! = 165 * 310* 8* 12*3 / 1.96785* 0.375

OpenStudy (anonymous):

well the choices are a)6.2x10^10 b)8.7x10^9 c)9.6x10^8 d)3.5x1^&10 e)2.9x10^11

OpenStudy (anonymous):

i make it 9.6 x 10^8

OpenStudy (anonymous):

958199862 to be exact

OpenStudy (anonymous):

jimmyrep is correct... I didn't account for inches vs. feet

OpenStudy (anonymous):

how did you get that? 165 * 310* 8* 12*3 / 1.96785* 0.375 why did you do *3?

OpenStudy (anonymous):

converting cu ft to cu inches - 1 cu ft = 12x12x12 = 12^3 cu ins

OpenStudy (anonymous):

oh alright i thought when you did LxWxH and the *12 it went from cuft to cuin

OpenStudy (anonymous):

i doubt if that lot would be very stable - especially in a gale !!!!

OpenStudy (anonymous):

no that just ft to ins

OpenStudy (anonymous):

lol its just a chem question but it would be pretty cool to see

OpenStudy (anonymous):

right

OpenStudy (anonymous):

wait would it be 12*12*12? or 12^3?

OpenStudy (anonymous):

lol nvm dumb question

OpenStudy (anonymous):

thats same thing 12^3 is a short way of saying 12x12x12 like 6^2 = 6x6

OpenStudy (anonymous):

the number after ^ is called the power or the exponent

OpenStudy (anonymous):

haha ok ty just its just been a long weekend for me

OpenStudy (anonymous):

i know the feeling - none of us are computers

OpenStudy (anonymous):

thats probably a good thing

OpenStudy (anonymous):

The problem can be restated as follows: How many match boxes, with the given dimension, can fit into a shipping container, measuring 165ft wide x 310ft long x 8 feet high? This is not a trivial problem in the real world. Check out the paper at: http://www.cs.nott.ac.uk/~sda/research/mipmodel.pdf